### v1

@@ 3.3.1 Retrieval Metrics
Non-Rank Based Metrics
Non-rank-based metrics often assess binary outcomes—whether an item is relevant or not—without considering the position of the item in a ranked list. Notice, that the following formula is just one format of these metrics, the definition of each metric may vary by the different evaluating tasks.
@@
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
    where $TP$ is the number of true positives, $TN$ is the number of true negatives, $FP$ is the number of false positives, and $FN$ is the number of false negatives.
    Precision is the fraction of relevant instances among the retrieved instances.
@@
    Recall at k ($Recall@k$) is the fraction of relevant instances that have been retrieved over the total amount of relevant instances, considering only the top $k$ results.
@@
Rank-Based Metrics
Rank-based metrics evaluate the order in which relevant items are presented, with higher importance placed on the positioning of relevant items at the top of the ranking list.
@@
    where $P(k)$ is the precision at cutoff $k$ in the list, $rel(k)$ is an indicator function equaling 1 if the item at rank $k$ is a relevant document, 0 otherwise, and $n$ is the number of retrieved documents.
@@ 3.3.2 Generation Metrics
    ROUGE Recall-Oriented Understudy for Gisting Evaluation (ROUGE) [32] is a set of metrics designed to evaluate the quality of summaries by comparing them to human-generated reference summaries. The variants of ROUGEs measure the overlap of n-grams (ROUGE-N, ROUGGE-W), word subsequences (ROUGE-L, ROUGGE-S), and word pairs between the system-generated summary and the reference summaries. ROUGE can be indicative of the content overlap between the generated text and the reference text.
    BLEU Bilingual Evaluation Understudy (BLEU) [36] is a metric for evaluating the quality of machine-translated text against one or more reference translations. BLEU calculates the precision of n-grams in the generated text compared to the reference text and then applies a brevity penalty to discourage overly short translations. BLEU has been applied to other generation tasks. Despite its popularity, BLEU has limitations, such as not accounting for the fluency or grammaticality of the generated text.
@@
    LLM as a Judge Using LLMs as judges for evaluating generated text is a more recent approach. [58] In this method, LLMs are used to score the generated text based on criteria such as coherence, relevance, and fluency. The LLM can be fine-tuned on human judgments to predict the quality of unseen text or used to generate evaluations in a zero-shot or few-shot setting. This approach leverages the LLM’s understanding of language and context to provide a more nuanced assessment of text quality. For instance, [1] illustrates how providing LLM judges with detailed scoring guidelines, such as a scale from 1 to 5, can standardize the evaluation process. This methodology encompasses critical aspects of content assessment, including coherence, relevance, fluency, coverage, diversity, and detail - both in the context of answer evaluation and query formulation.
@@ 3.3.3 Additional Requirements
Latency
measures the time taken by the RAG system to retrieve relevant information and generate a response. It is a critical factor for user experience, especially in interactive applications such as chatbots or search engines [18].
Single Query Latency: The average time taken to process a single query, including both retrieval and generating phases.
Diversity
evaluates the variety and breadth of information retrieved and generated by the RAG system. It ensures that the system can provide a wide range of perspectives and avoid redundancy in responses [4].
Cosine Similarity / Cosine Distance: The cosine similarity/distance calculates embeddings of retrieved documents or generated responses. Lower cosine similarity scores indicate higher diversity, suggesting that the system can retrieve or generate a broader spectrum of information.
Noise Robustness
measures the RAG system’s ability to handle irrelevant or misleading information without compromising the quality of the response [33]. The metrics Misleading Rate and Mistake Reappearance Rate are described in [33], providing detailed descriptions tailored to the specific dataset and experimental setup.
Negative Rejection
evaluates the system’s capability to withhold responses when the available information is insufficient or too ambiguous to provide accurate answer [5].
Rejection Rate: The rate at which the system refrains from generating a response.
Counterfactual Robustness
Counterfactual robustness assesses the system’s ability to identify and disregard incorrect or counterfactual information within the retrieved documents [34].
Error Detection Rate: The ratio of counterfactual statements are detected in retrieved information.
@@ 4 Discussion
Evaluating Retrieval-Augmented Generation (RAG) systems comprises multifaceted challenges, arising from their dual reliance on retrieving accurate, relevant information and generating coherent responses that meet user expectations. This survey has highlighted several key points of consideration, capturing the breadth and depth of evaluations necessary for advancing RAG technologies.
Within the evaluation target aspect, it’s evident that traditional question-answering (QA) setups remain effective for evaluating generative components, particularly in how well they align with human preferences for clarity, relevance, and factual accuracy. However, recent strategies, such as CRUD-based assessments, offer novel angles by scrutinizing RAG systems’ interactive capabilities with dynamic information environments [34]. These methodologies underscore the necessity for RAG evaluations to evolve beyond static benchmarks, mirroring real-world scenarios where information is continuously updated and queries are not strictly fact-based but exploratory or conversational.
On the dataset front, the challenge of devising a "one-size-fits-all" dataset is pronounced, given the highly task-specific nature of RAG systems. Unique datasets, meticulously crafted to test specific facets of RAG performance, are indispensable. While this approach ensures thorough, targeted evaluation, it also magnifies the effort and resources required for comprehensive testing. The divergence of datasets, ranging from news articles to structured databases, reflects the adaptability required of RAG systems but also signifies a formidable barrier to streamlined evaluation [49, 50].
@@
When it comes to metrics, the use of LLMs as automatic evaluative judges signifies a burgeoning trend, promising versatility and depth in generative outputs with reasoning on a large scale compared to human evaluation. However, using LLMs as judges for chatbot responses presents challenges in aligning with human judgment, establishing effective grading scales, and applying consistent evaluation across varied use cases. The determination of correctness, clarity, and richness can differ between automated and human assessments. Moreover, the effectiveness of example-based scoring can vary, and there’s no universally applicable grading scale, complicating the standardization of LLM as a judge. [28]
Future directions in RAG evaluation should focus on developing more adaptive, context-aware benchmarks that accurately reflect the dynamic, information-rich environments these systems are designed to navigate. Such efforts could include simulating real-time information updates in evaluation datasets or incorporating user feedback loops into assessment methodologies. Additionally, exploring more nuanced metrics that can capture the subtlety of human language comprehension and generation—beyond sheer accuracy or relevance—will be crucial. Efforts to codify these advancements into standardized evaluation frameworks would significantly bolster the field, providing clearer benchmarks for progress and more directly aligning RAG system advancements with user needs and societal impacts.
@@ 5 Conclusion
This survey has systematically explored the complex landscape of evaluating Retrieval-Augmented Generation (RAG) systems, highlighting the multifaceted challenges inherent in assessing their performance. Through the proposed RGAR analysis framework, we have delineated a structured approach to dissecting the intricacies of RAG evaluations, focusing on the retrieval, generation, and additional requirements that underpin these systems. Our comprehensive analysis underscores the necessity for targeted benchmarks that reflect the dynamic interplay between retrieval accuracy and generative quality, as well as the practical considerations crucial for real-world applications. By identifying gaps in current methodologies and suggesting future research directions, this survey aims to pave the way for more nuanced, effective, and user-aligned evaluations of RAG systems, ultimately contributing to the advancement of natural language processing technologies that are both intelligent and versatile.

### v2

@@ 3.3.1 Retrieval Metrics
Non-Rank Based Metrics often assess binary outcomes—whether an item is relevant or not—without considering the position of the item in a ranked list. Notice, that the following formula is just one format of these metrics, the definition of each metric may vary by the different evaluating tasks.
@@
    Precision is the fraction of relevant instances among the retrieved instances,
@@
    Recall at k ($Recall@k$) is the fraction of relevant instances that have been retrieved over the total amount of relevant cases, considering only the top $k$ results.
@@
Rank-Based Metrics evaluate the order in which relevant items are presented, with higher importance placed on the positioning of relevant items at the ranking list.
@@
    where $P(k)$ is the precision at cutoff $k$ in the list, $rel(k)$ is an indicator function equaling 1 if the item at rank $k$ is a relevant document, 00 otherwise, and $n$ is the number of retrieved documents.
@@ 3.3.2 Generation Metrics
    ROUGE Recall-Oriented Understudy for Gisting Evaluation (ROUGE) [37] is a set of metrics designed to evaluate the quality of summaries by comparing them to human-generated reference summaries. ROUGE can be indicative of the content overlap between the generated text and the reference text. The variants of ROUGEs measure the overlap of n-grams (ROUGE-N, ROUGGE-W), word subsequences (ROUGE-L, ROUGGE-S), and word pairs between the system-generated summary and the reference summaries.
    BLEU Bilingual Evaluation Understudy (BLEU) [43] is a metric for evaluating the quality of machine-translated text against one or more reference translations. BLEU calculates the precision of n-grams in the generated text compared to the reference text and then applies a brevity penalty to discourage overly short translations. BLEU has limitations, such as not accounting for the fluency or grammaticality of the generated text.
@@
    LLM as a Judge Using “LLM as a Judge” for evaluating generated text is a more recent approach. [75] In this method, LLMs are used to score the generated text based on criteria such as coherence, relevance, and fluency. The LLM can be optionally finetuned on human judgments to predict the quality of unseen text or used to generate evaluations in a zero-shot or few-shot setting. This approach leverages the LLM’s understanding of language and context to provide a more nuanced text quality assessment. For instance, [1] illustrates how providing LLM judges with detailed scoring guidelines, such as a scale from 1 to 5, can standardize the evaluation process. This methodology encompasses critical aspects of content assessment, including coherence, relevance, fluency, coverage, diversity, and detail - both in the context of answer evaluation and query formulation.
@@ 3.3.3 Additional Requirements
Latency measures the time taken by the RAG system to finish the response of one query. It is a critical factor for user experience, especially in interactive applications such as chatbots or search engines [20]. Single Query Latency: The mean time is taken to process a single query, including both retrieval and generating phases.
Diversity evaluates the variety and breadth of information retrieved and generated by the RAG system. It ensures that the system can provide a wide range of perspectives and avoid redundancy in responses [4]. Cosine Similarity / Cosine Distance: The cosine similarity/distance calculates embeddings of retrieved documents or generated responses. [30] Lower cosine similarity scores indicate higher diversity, suggesting that the system can retrieve or generate a broader spectrum of information.
Noise Robustness measures the RAG system’s ability to handle irrelevant or misleading information without compromising the quality of the response [38]. The metrics Misleading Rate and Mistake Reappearance Rate are described in [38], providing detailed descriptions tailored to the specific dataset and experimental setup. [58]
Negative Rejection evaluates the system’s capability to withhold responses when the available information is insufficient or too ambiguous to provide an accurate answer [6]. Rejection Rate: The rate at which the system refrains from generating a response.
Counterfactual Robustness Counterfactual robustness assesses the system’s ability to identify and disregard incorrect or counterfactual information within the retrieved documents [39]. Error Detection Rate: The ratio of counterfactual statements detected in retrieved information.
@@ 4 Discussion
For RAG systems, traditional Question Answering (QA) datasets and metrics remain a common format for interaction. [14, 49, 38, 6, 61, 62, 58, 66] While these provide a basic verification of RAG’s capabilities, it becomes challenging to distinguish the impact of retrieval components when faced with strong Language Models (LLMs) capable of excelling in QA benchmarks. To comprehensively evaluate the performance of entire RAG systems, there is a need for diverse and RAG-specific benchmarks. Several papers offer guidance on improving QA format benchmarks, including variations in question types: from simple Wikipedia filling questions to multi-hop [52], multi-document questions [66] and single-round to multi-round dialogue [39, 58]. For answers, aspects such as structural output [58], content moderation [6, 54], and hallucination [66] can be considered when evaluating relevance, faithfulness, and correctness. In addition to these, RAG systems require additional requirements such as robustness to noisy documents, language expression, latency, and result diversity. [32, 33, 38, 6, 39, 57, 58, 20, 4] Furthermore, research is needed on performance changes involving intermediate outputs and retrieved documents, as well as the relationship and analysis between retrieval metrics and final generation outputs.
Regarding datasets, creating a universal dataset was challenging due to the target-specific nature of different RAG benchmarks. Tailored datasets [14, 38, 49, 39, 57] are necessary for a thorough evaluation, but this approach increases the effort and resources required. Moreover, the diversity of datasets, from news articles to structured databases [66], reflects the adaptability required of RAG systems but also poses a barrier to streamlined evaluation. Recently, with the cutting-edge performance of LLMs, complex data processing and automatic QA pair generation can be automated to achieve daily or finer-grained time resolution, preventing LLMs from cheating and evaluating the robustness of RAG systems in rapidly changing data. [6, 52, 39, 62, 58, 66]
@@
When it comes to metrics, the use of LLMs as automatic evaluative judges signifies a burgeoning trend, promising versatility and depth in generative outputs with reasoning on a large scale compared to human evaluation. However, using “LLMs as a Judge” [75] for responses presents challenges in aligning with human judgment, establishing effective grading scales, and applying consistent evaluation across varied use cases. Determining correctness, clarity, and richness can differ between automated and human assessments. Moreover, the effectiveness of example-based scoring can vary, and there’s no universally applicable grading scale and prompting text, complicating the standardization of “LLM as a Judge”. [33]
In addition to the challenges mentioned above, it is important to consider the resource-intensive nature [76] of using Large Language Models (LLMs) for data generation and validation. RAG benchmarks must balance the need for thorough evaluation with the practical constraints of limited computational resources. As such, it is desirable to develop evaluation methodologies that can effectively assess RAG systems using smaller amounts of data while maintaining the validity and reliability of the results.
@@ 5 Conclusion
This survey systematically explores the complexities of evaluating RAG systems, highlighting the challenges in assessing their performance. Through the proposed A Unified Evaluation Process of RAG, we outline a structured approach to analyzing RAG evaluations, focusing on targets, datasets and measures. Our analysis emphasizes the need for targeted benchmarks that reflect the dynamic interplay between retrieval accuracy and generative quality and practical considerations for real-world applications. By identifying gaps in current methodologies and suggesting future research directions, we aim to contribute to more effective, and user-aligned benchmarks of RAG systems.